8. Given the function g(x) = 0.75 cos(x – 0.5, the period is a. 4 b. EIN c. d. 0.5
Answer #1

Option a is correct. Explanation are given below

solutions et be the Period of gens then, gentas -0.5 = 0.75 ces centa) = 0.75 (cas (An (cas (n + 13)) 0.5 = 0.75 (cas (5) eos